Sol Summary – December 3rd

Crew 184 Daily Summary Report  03 Dec 2017

MDRS Daily Summary Report for sol 2

Summary Title:  Getting settled at MDRS!

Mission Status:  Today has been a busy day preparing both ourselves and the MDRS for the rest of our mission.  We spoiled ourselves with a hearty breakfast today (and to get rid of some of our fresh food!).

Today has been one of excitement preparing for our first EVA tomorrow.  We have encountered our first mission failure!

During the high winds today we heard a bang as the outer airlock door was blown open.  The inner airlock door held closed and thus we didn’t all die!  We immediately closed the door to the EVA Prep room to form a second seal to the martian atmosphere.  During tomorrows familiarization EVA we will enter the EVA Prep room, close the EVA Prep room door, and use that as a depressurization chamber.  Immediately upon departure on tomorrows EVA we will repair the outer airlock door.

We have discovered today that all the planning in the world still requires real-time flexibility as numerous activities took much longer than we expected.  For example the Exercise Bike took several hours and modifications to assemble and get working correctly, much more than we had allocated.  That being said, we have successfully completed all objectives and are ready to begin completing our science objectives tomorrow!.

Sol Activity Summary:

1.    Inventoried and organized the MDRS Hab and Greenhab

2.    Completed all reports

a.    Greenhab

b.    HSO

c.    Mission Plan

d.    Operations Report

e.    Sol Summary

f.     EVA Request

g.    Journalist

3.    Assembled exercise equipment (Stationary Bike / Weights)

4.    Got more detailed training from previous crewmembers

5.    Crew Exercise

6.    Internet troubleshooting (WiFi modem nonfunctional – connection working straight to modem)

7.    Cleaning and organizing MDRS

8.    Radio check with all radios (all functional except for 1)

9.    Communication check at all MDRS buildings with Habitat

10. Neuroblock Feasability Training

Look Ahead Plan:

1.    Familiarization EVA

2.    Communication ‘Lego Bricks’ experiment

3.    Sleep Study / crew medical questionnaires

4.    Musk Observatory Training and Observation

5.    Journalist interviews and filming

6.    Medical status surveys with Crew Doctor throughout the day.

7.    CPR Training

a.    Practice CPR techniques with crew medical doctor on CPR medical dummy

See schedule for tomorrow attached to this email.

Anomalies in work:

8.    Non Functional WiFi Modem

9.    Non Functional Power Strip (removed and replaced)

10. 1 Non Functional Radio

11. EVA Airlock Outer Hatch Failure (failed open) due to high winds.

12. Roof Hatch blown open continually (will close when winds die down)

Weather:  Very windy!  We have had two wind related MDRS failures.

Crew Physical Status:  All crew in good health.

EVA: No EVA today

Reports to be filed:

Greenhab

Operations

EVA Request

Journalist

Sol Summary

Mission Plan

Support Requested:

1.    We can’t find a template for Daily HSO Report.  Please provide what is desired in this report.

2.    Water Transfer from Delivery Tank to Static Tank

3.    More TB-Cide Quat Cleaner (we are out)

4.    Recommendation for wi-fi router failure.

a.    We have reset the router, checked our allocation, restarted laptop, and router is non-functional.  When hooking up laptop directly to modem internet works well.

Greenhab Report – December 3rd

GreenHab Report

Trisha Randazzo

December 3, 2017

Sol 2

Environmental control: (Choose which is appropriate and explain further if needed)

Ambient (no heating or/cooling)

Ambient with window/door open

· Windy so unable to keep door open

Heating

· Functioning nominally

Cooling

· Not availabloe

Both heating and cooling

· Nominal

Shade cloth on/off

· On

Average temperature: (N/A until new sensor is delivered)

· N/A

Changes to crops: Note all emergence of seedlings, death of seedlings, etc

· One tomato is doing poorly, but everyone else is growing quite well!!

Daily water usage for crops:

· Complete with no issues

Time(s) of watering for crops:

· 12:55

Morning research observations: (Currently not operational)

· N/A

Changes to research plants:

· N/A

Daily watering and amount of water used:

· 2 gallons of all soils

Aquaponics: (Currently not operational)

· N/A

Narrative: Any other information you want to share

· Planted crew gifts of paper-whites. Two containers in the Green-Hab, one container in the Hab.

Support/supplies needed:

· Awaiting:

Sol Summary – December 2nd

Sol 1

Summary: MDRS Training Day!

Author: Tom Horn

Mission Status:  Mission is starting great and we are enjoying our first day on Mars! Today was a busy day as we spent the day out of sim getting training from the previous crew and Shannon. We are excited for our first day ‘in sim’ tomorrow. We are taking a ‘light duty’ day tomorrow as we focus on finishing move in to the hab and getting acclimated. Due to this we are NOT doing an EVA tomorrow.

Notes for CAPCOM. Please communicate with our crew using this email address, mdrscrew184@gmail.com. This ensures our entire crew sees all communications. Please add this email address to the mission support distribution so in the future we can email these reports to you from that address.

One of the major research goals for our team is to follow a ‘Martian Day’ of 24:40 and study the effects on our team from following a longer day. Our team is going to follow the schedule below. We understand the the ‘CAPCOM Communication Window’ is between 7-9pm, we will always be awake during this time and will ensure to get reporting completed during this window, including any and all EVA requests. Note that we won’t be following a normal day so if a request is sent to our crew during our ‘sleep time’ our response will be delayed.

Further details will follow in the Mission Plan to be sent tomorrow.

Flight Day Calendar Day Wake Bed Twilight Start Sunrise Sunset Twilight End
Day 1 12/2 Not Scheduled 8:00 PM 6:53:57 AM 7:23:08 AM 5:02:24 PM 5:31:34 PM
Day 2 12/3 4:00 AM 8:40 PM 6:54:50 AM 7:24:03 AM 5:02:17 PM 5:31:30 PM
Day 3 12/4 4:40 AM 9:20 PM 6:55:41 AM 7:24:57 AM 5:02:13 PM 5:31:28 PM
Day 4 12/5 5:20 AM 10:00 PM 6:56:32 AM 7:25:50 AM 5:02:11 PM 5:31:28 PM
Day 5 12/6 6:00 AM 10:40 PM 6:57:21 AM 7:26:41 AM 5:02:10 PM 5:31:30 PM
Day 6 12/7 6:40 AM 11:20 PM 6:58:10 AM 7:27:32 AM 5:02:12 PM 5:31:34 PM
Day 7 12/8 7:20 AM 12:00 AM 6:58:57 AM 7:28:22 AM 5:02:16 PM 5:31:41 PM
Day 8 12/9 8:00 AM 12:40 AM 6:59:44 AM 7:29:10 AM 5:02:23 PM 5:31:49 PM
Day 9 12/10 8:40 AM 1:20 AM 7:00:29 AM 7:29:56 AM 5:02:31 PM 5:31:59 PM
Day 10 12/11 9:20 AM 2:00 AM 7:01:13 AM 7:30:42 AM 5:02:42 PM 5:32:11 PM
Day 11 12/12 10:00 AM 2:40 AM 7:01:55 AM 7:31:26 AM 5:02:54 PM 5:32:25 PM
Day 12 12/13 10:40 AM 3:20 AM 7:02:36 AM 7:32:08 AM 5:03:09 PM 5:32:41 PM
Day 13 12/14 11:20 AM 4:00 AM 7:03:16 AM 7:32:49 AM 5:03:26 PM 5:32:58 PM
Day 14 12/15 12:00 PM 4:40 AM 7:03:55 AM 7:33:29 AM 5:03:44 PM 5:33:18 PM
Day 15 12/16 12:40 PM Not Scheduled 7:04:32 AM 7:34:06 AM 5:04:05 PM 5:33:40 PM
Day 16 12/17 Not Scheduled – Handover Day 7:05:07 AM 7:34:42 AM 5:04:28 PM 5:34:03 PM

 

Sol Activity Summary: We arrived at 12pm today and spent the rest of the day learning about the MDRS with Shannon. We enjoyed our last ‘out of sim’ dinner at Stan’s burger shack with Shannon and the previous crew.

Look Ahead Plan:

Tomorrow our primary tasks are:

  1. Finishing move in to the HAB
  2. Completing an MDRS Inventory of supplies (food and other)
  3. Finish cleaning MDRS to start the mission.
  4. ScienceDome / Greenhab / Solar Observatory Detailed Inspection
  5. Completion of required reporting
    1. Mission Plan
    2. Health and Safety Officer Report
    3. If other ‘beginning of sim’ reports are requested, please let us know!

Anomalies in work:  None

Weather: Unseasonably warm!

Crew Physical Status: All crew in good physical health. Brandon Ferguson from the previous crew is staying with our crew until Monday night.

EVA: No EVA

Reports to be filed:

  1. Sol Summary (this is intended to include the commanders report as a narrative included in Mission Status. If this did not hit sufficient detail for this please let me know what details are desired and I will include them in future reports)
  2. Operational Report

Support Requested: None

Journalist Report & Crew Photos – December 2nd

Crew 184 Journalist Report

Willie Schumann

02 December 2017

Title Day One on Mars

Narrative We are on Mars. It’s unbelievable, but yet so real. After half year of planning , we arrived yesterday to the habitat and it was a rush of first impressions. We cleaned some larger items of the previous crews and made ourselves at home. The effects of the long Martian day, which is 40 minutes longer then a day on earth, kicked in and we were quite tired. We dropped into our beds happy and exhausted, which is the best state of mind when you finish a day.

There is always beauty in new beginnings, and so was the morning of our first Martian Day. We got up really early. There is still so much to be done before we can start our research and breath life into the plentiful projects of crew 184. Nonetheless, we spoiled ourselves with a hearty breakfast of scrambled eggs. We are trying to use our fresh food as soon as possible. Especially in the first days we want to keep moral high, as the six of us have to get used to the new environment.

After breakfast, our second in command Trisha Randazzo gave everyone in the crew a little bag of presents. What a surprise, which could only be topped by Mars itself. The first sunrise on the red planet blew our minds. The colors of the distant sun, leaping over the Martian hills was something I had never seen on Earth. Maybe the euphoria tainted my perception, but anyway, it was darn beautiful.

Even though we were not assigned on an EVA, just yet, the day was eventful. First Officer Randazzo and Crew Engineer Hunt started assembling the workout bicycle really early and had a steep learning curve. It’s seems to be a universal truth, that assembling is as hard on Mars as it is on Earth, and that improvisation is always king. It’s not important how you get there, but that is works and that you gained knowledge in the process.

It’s not a surprise that the crew meals on this first day have been a very bonding experience. I hope we can keep this energy, even if we run out of fresh food and of inspiration how to prepare it. We cherish being here together and looking forward for the days to come. We planted six paper white flower in the green hab this noon, one for every crewmember. So we set already the first seeds of new life on this strange planet. We try to keep them alive. If we succeed, we will probably manage most other challenges that wait for us.

Tomorrow we will head out for our first EVA’s. I think, that is when the magic really starts. We are familiar with our new home now, but now we want to see the land, that surrounds us. We hope we can manage to wear the new space suits without larger problems, as we have heard, that you have to be fit. That’s another reason, why we plan an intense workout regime during our stay. I think, even if we struggle in the beginning of our mission, practice will show its effect and we can start to enjoy our exploration of Mars to the fullest.

Personal Logbook As a Journalist and filmmaker I have certain goals like everyone else of my crewmates. I am very excited about the time to come and conscious about my tasks and the things I need to do, to create great stories. While today was mainly spent with preparations and adjustments, tomorrow will be opening a great window of opportunity for fantastic pictures, when we will finally go outside.

I have filmed a lot today, but I think, that the scenes, that follow, will be more meaningful. As now, I haven’t managed to make to many interviews and it will be a challenge to find new corners and angles in the habitat to film. I was a little tired this morning, but worked on autopilot through most of the day. Patience and motivation will be important to make the most of my time on Mars. I guess, these are qualities that are essential for every astronaut and storyteller.

Willie Schumann, Journalist, Crew 184

Operations Report – December 2nd

Crew 184 Operations Report 02/12/2017

SOL: SOL 1

Name of person filing report: HuntJ

Non-nominal systems: All Nominal

Notes on non-nominal systems: N/A

Generator (hours run): N/A (Arrived during day when already off)

Solar— SOC % (Before generator is run at night) 80%

Diesel – 50% Full.

Propane – 80 Percent Volume .

Ethanol Free Gasoline (5 Gallon containers for ATV) – gallons. 5 gallons

Water (trailer) – 1000 Gallons!!!.

Water (static) – 500 gallons

Trailer to Static Pump used – No

Water (loft) – Static to Loft Pump used – No

Water Meter: 46 gallons

Toilet tank emptied: No

ATV’s Used: Honda, 300, 350.1, 350.2, 350.3

Oil Added? No

ATV Fuel Used: Negligible (Test drive only)

# Hours the ATVs were Used today: 00:20 hours

Notes on ATVs: They’re running great!!

Deimos rover used: Not used

Hours:

Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: Not used

Hours:

Beginning charge:

Ending charge:

Currently charging:

Opportunity rover used: Not Used

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used: Not Used

Hours:

Beginning charge:

Ending charge:

Currently charging:

HabCar used and why, where? No

General notes and comments: Happy to see the Red Planet!

Summary of internet: So far so good!

Summary of suits and radios: All nominal

Summary of Hab operations: All nominal

Summary of GreenHab operations: All nominal

Summary of ScienceDome operations: All nominal

Summary of RAMM operations: Not Operational

Summary of health and safety issues: All nominal

Questions, concerns and requests to Mission Support: Everything is Awesome!!

Sol Summary – December 1st

Crew 183+ Sol Summary Report 1 December, 2017

Sol 4

Title: Curtis Arrives!

Written by: Brandon Ferguson

Mission Status: Good. Operations are Normal

Sol Activity Summary:

Today, we spent most of our time at the hab. At long last, Curtis arrives to fix the water heater in the hab, which has been out of commission for over a month now. He also insulated the water pipes and the outside tank. This morning, I was able to start my final batch of Sedimentation tests. All that’s left are my salinity & pH tests for all the soils.

Meanwhile, Chris offered to help put the rovers together. We now have three out of the four new rovers fully assembled.

At about 4:00 PM, I took a short EVA to Tank wash to resupply on a couple of samples that I needed.

Look Ahead Plan:

Tomorrow morning, Curtis will deliver and install Shannon’s water tank tomorrow, Shannon and I will be preparing the Lab and the Hab for Crew 184’s arrival

Anomalies in work: No Problems

Weather: Clear, warm, calm winds

Crew Physical Status: OK, Healthy

EVA: Tank Wash area

Reports to be Filed:

-Operations Report

Support Requested: N/A

Operations Report – December 1st

Crew 183 + Operations Report 1 Dec 2017

SOL: n/a

Name of person filing report: Shannon Rupert

Non-nominal systems: Phone is still not functional, Low water pressure in Hab

Notes on non-nominal systems: HudgesNet was here all day but unable
to repair the phone. They will be working on it again on Monday. The
filter than normally clogs up the water line in the Hab is missing,
but two other filters were cleaned (Bathroom and kitchen sink) and
water pressure is once again nominal throughout the Hab. A new master
filter will be installed to forever end this problem, which was
probably why the old water heater died.

Generator (hours run): off 8 am on 6 pm

Solar— SOC 94% (Before generator is run at night)

Diesel – 50 %

Propane – 100 psi

Ethanol Free Gasoline (5 Gallon containers for ATV) – not known gallons.

Water (trailer) – 0 gallons. Curtis took the water trailer to town to fill it.

Water (static) – not known gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: not known

Toilet tank emptied: no

ATV’s Used: yes

Oil Added? No

ATV Fuel Used: 0 Gals

# Hours the ATVs were Used today: 2

Notes on ATVs: nothing to report

Deimos rover used: yes

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: no

Sojourner rover used: ASSIGNED TO DIRECTOR, Operational

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: Operational

Hours:

Beginning charge:

Ending charge:

Currently charging:

Opportunity rover used: Operational

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used: Not operational

Hours:

Beginning charge:

Ending charge:

Currently charging:

HabCar used and why, where? No

General notes and comments: Remember that for the first month of use,
the new rovers will not hold a charge as well as Deimos. We took the
three new ones on a 5 mile round trip and they were each reduced to
about 75%. It takes a while for the batteries to perform at their
peak.

I’d like to say that the rovers were made operational by Brandon and I
but it was actually Chris who did a majority of the work and taught me
a lot about the right tools for the job. Who is Chris? Well, he’s
the guy who was here to fix my phone. It took so long he offered to
help.

The static tank was insulated today. A new heater was installed and
the pipe to the Hab was insulated as well. Shannon’s water tank was
delivered and will be installed tomorrow.

Water heater: After determining that the new water heater would not
fit in the spot where the old one was mounted, it was moved to a spot
on the lower deck just to the right of the ductless propane heater.
It was operational, but would not function correctly until the water
pressure was restored. It is currently connected to the heater’s
propane line. I’m assuming Curtis will reconnect the heater as soon
as he has the supplies.

Summary of internet: not known

Summary of suits and radios: nothing to report

Summary of Hab operations: Water system insulated on the Hab’s
exterior. New hot water heater installed in new location. Water
pressure restored. Water trailer to be refilled.

Summary of GreenHab operations: The thinned plants look mostly good.
One of the tomatoes I tried to transplant does not look good. The
rest look fine. Door was open for a few hours today but it was cold
enough that it stayed mostly warm and not hot.

Summary of ScienceDome operations: Brandon concluded his work.
Nothing more to report.

Summary of RAMM operations: nothing to report

Summary of health and safety issues: nothing to report

Questions, concerns and requests to Mission Support: Nothing, thank you.

Sol Summary – November 30th

Crew 183+ Sol Summary Report 30November2017

Sol 3

Title: Further Exploration

Written by: Brandon Ferguson

Mission Status: Good. Operations are Normal

Sol Activity Summary:

Today, I spent from 10:00 AM to 2:00 PM, exploring the area south of Kissing Camel Range. I’d estimate that I covered anywhere from 6-8 km during that time frame. There were many interesting areas that I found, such as a couple of caves that a person can walk into. Although I didn’t find any dinosaur bones out there, I did find a lot of petrified wood. During this time, Shannon managed to complete the MDRS user manual that she’s been working on.

Between 3:00 PM & 5:00 PM, Shannon and I explored the last of the unchecked roads we still needed to check up on.

Look Ahead Plan:

Tomorrow, Curtis and his team will come and fix the water heater. Shannon and I will clean up the Lab for Crew 184.

Anomalies in work: No Problems

Weather: Clear, warm, calm winds

Crew Physical Status: OK, Healthy

EVA: Region South of Kissing Camel Range & Road exploration

Reports to be Filed:

-Operations Report

Support Requested: N/A

Operations Report – November 30th

Crew 183 + Operations Report 30NOV2017

SOL: n/a

Name of person filing report:  Shannon Rupert

Non-nominal systems: Nothing new to report

Notes on non-nominal systems:  n/a

Generator (hours run): off 8 am on 6 pm

Solar— SOC 90% (Before generator is run at night)

Diesel –  50 %

Propane – 100 psi

Ethanol Free Gasoline (5 Gallon containers for ATV) – not known gallons.

Water (trailer) – 0 gallons.

Water (static) – 250 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter:  not known

Toilet tank emptied: no

ATV’s Used: Honda

Oil Added? No

ATV Fuel Used:  2 Gals

# Hours the ATVs were Used today: 4

Notes on ATVs: nothing to report

Deimos rover used:  yes

Hours:

Beginning charge: 100

Ending charge: 94

Currently charging: yes

Sojourner rover used:  ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used:  Installation in progress

Hours:

Beginning charge:

Ending charge:

Currently charging:

Opportunity rover used:  Installation in progress

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used:

Hours:

Beginning charge:

Ending charge:

Currently charging:

HabCar used and why, where?  No

General notes and comments: Burn barrel not complete. Turns out you
can’t just poke holes in it with a pickaxe.  We have a plan for
tomorrow.  Also, contractors will be onsite all day tomorrow.

Summary of internet:  not known

Summary of suits and radios: nothing to report

Summary of Hab operations: nothing to report

Summary of GreenHab operations: Door was opened this afternoon.
Thinned the babies.  Used 4 gallons of water for watering and prepping
the soil for herbs.

Summary of ScienceDome operations:  Brandon continues his work.
Nothing more to report.

Summary of RAMM operations: nothing to report

Summary of health and safety issues: nothing to report

Questions, concerns and requests to Mission Support: Nothing tonight, thank you.

Sol Summary – November 29th

Sol 2

Title: An Intriguing Find…

Written by: Brandon Ferguson

Mission Status: Good. Operations are Normal

Sol Activity Summary:

Today, I set off to explore the area to the South of Kissing Camel Range and West of Cow Dung Rd. First I investigated some of the washes crossing the road to see what I could find. In the first wash that I stopped in, I found pieces of petrified wood amongst other very interesting specimens. I’ll definitely have to explore this drainage basin later.

From there, I proceeded further south, until I reached a large wash near White Rock Canyon. I only walked about 2 minutes up, when I found a very large specimen that looked like mineralized bone marrow. Compared to the basalt ejecta littering the wash, this specimen looked completely different. The holes were all aligned in the same direction, and were about the same size. The basalt pieces had gas bubbles, but were of random size, with no obvious pattern. The size of the marrow was very large, with 3-4 mm spacing. If this was indeed a piece of a dinosaur bone, then that meant there is a very large dinosaur fossil somewhere in that drainage basin.

I left the specimen where I found it, planning to come back and visit later. After spending over 2 hours in this area, I didn’t find much more, though there is still quite a bit of ground to cover. I returned to the hab by around 2:00 PM, and informed Shannon on what I had found. On our way to town, we investigated the specimen that I had found in the wash. Shannon wasn’t sure what it was, but we confirmed that there weren’t any other rocks in that area that looked like it.  She figured that if it was indeed a remnant of a dinosaur bone, that it would have to come from a very large dinosaur, much larger than a Utahraptor or velociraptor. We plan on surveying this area further to see what we can find.

Soon after, we ran into town, found a socket wrench kit that would work with the bolts on the new rovers, and organized the next supply of food for Crew 184. We returned to hab after finishing our business in town.

Look Ahead Plan:

Tomorrow and the day after, I will be extensively searching the drainage basin where I found that specimen for large fossil exposures. Since Shannon told me that she hasn’t extensively explored this area, what I may find there could be very interesting.

Anomalies in work: No Problems

Weather: Clear, warm, calm winds

Crew Physical Status: OK, Healthy

EVA:

Reports to be Filed:

-Operations Report

Support Requested: N/A